the base of an isoceles triangle is 10cm and one of its sides is 8cm find the area of the triangle.

Answer by stanbon(75887)

the base of an isoceles triangle is 10cm and one of its sides is 8cm find the area of the triangle.
----
Draw an isoceles triangle.
Label the base "10".
Label one of the sides 8
---
Label the other side 8 because the sides are equal.
--
Draw an altitude down from the vertex to the base.
It is a perpendicaular-bisector of the base.
-----
Each section of the base is now "5".
---
Use Pythagoras to find the height:
8^2 = 5^2 + h^2
h = sqrt[64-25] = sqrt(39)
-------
Area = (1/2)(base)(height)
= (1/2)(10)*sqrt(39) is approximately 31.22 sq cm
===================================================
Cheers,
Stan H.
===================